PLEASE PLEASE PLEASE HELP ME! IT'S DUE IN 5 HOURS ONLINE! THANKS!
A right triangle has one leg three times as long as the other. Find a function that models its perimeter P in terms of the length x of the shorter leg.
P=

Answer by Alan3354(69443)

A right triangle has one leg three times as long as the other. Find a function that models its perimeter P in terms of the length x of the shorter leg.
P=
-------------------
hyp^2 = x^2 + (3x)^2
= 10x^2
The 3 sides are x, 3x and x*sqrt(10)
P = x + 3x + x*sqrt(10)
P = x*(4 + sqrt(10)), apx 7.162x